-1/4(x-12)+1/2(x=2)-x+4
Answer by stanbon(75887)

-1/4(x-12)+1/2(x+2)=x+4
---
Multiply thru by 4 to get:
----
-(x-12) + 2(x+2) = 4x+16
-----
-x+12 + 2x + 4 = 4x + 16
x + 16 = 4x + 16
3x = 0
x = 0
====================
Cheers,
Stan H.
========================
Cheers,
Stan H.